Solve the differential equation dy/dx+xy=y^2e^(x^2/2)sinx
1
Expert's answer
2020-03-21T17:12:07-0400

"y'+xy=y^2e^{x^2\/2}sinx"

divide both sides by "y^2"

"y'\/y^2+x\/y=e^{x^2\/2}sinx"

substitution: "t=1\/y"

"xt-t'=e^{x^2\/2}sinx"

let's solve first this equation: "xt-t'=0"

"\int dt/t=\int xdx"

"lnt-lnC=x^2\/2"

"t=Ce^{x^2\/2}"

then solution to initial equation will have form: "t(x)=C(x)e^{x^2\/2}"

"t'(x)=C'(x)e^{x^2\/2}+xC(x)e^{x^2\/2}"

Let's put these to initial equation

"xC(x)e^{x^2\/2}-C'(x)e^{x^2\/2}-xC(x)e^{x^2\/2}=e^{x^2\/2}sinx"

"C'(x)=-sinx"

"C(x)=\int -sinxdx=C_1+cosx"

then "t(x)=(C_1+cosx)e^{x^2\/2}"

"y=1\/t=e^{-x^2\/2}\/(C_1+cosx)"